LunaBechtold:

why does 2+2 equal fish and not 2+ backwards 2

Reeclark:

huh?

Midnight97:

it this a joke or something?

LunaBechtold:

midnight it is a question related to math

doggofroyo:

2+2=4

Midnight97:

@lunabechtold wrote:
midnight it is a question related to math
ik but i'm asking a question

LunaBechtold:

but i mean y doesnt 2+backwards 2 = fish

hannahj:

2 + 2 = 22 whatcha talking about

Midnight97:

2+2=4

LunaBechtold:

no 1+1=WiNdoW lolz

Midnight97:

but what i am trying to know is this a joke?

LunaBechtold:

no im serious bc idk why

LunaBechtold:

WHY ISNT IT THE SECOND ONE

Midnight97:

hmmm

Midnight97:

sorry this isn't math

LunaBechtold:

it kinda is tho

LunaBechtold:

ill close it

hannahj:

@lunabechtold wrote:
it kinda is tho
it kind of isnt tho

LunaBechtold:

i closed it

Midnight97:

@lunabechtold wrote:
ill close it
thank you and sorry

Arizona:

cause it =s 4 and fish has 4 letters